How many grams of NaN3 are required to form 1.00 g N2?


1
Expert's answer
2021-03-22T06:34:19-0400

Solution:

The balanced chemical equation:

2NaN3(s) → 2Na(s) + 3N2(g)

According to the equation above: n(NaN3)/2 = n(N2)/3


Moles of N2 = Mass of N2 / Molar mass of N2

The molar mass of N2 is 28.0134 g mol-1.

Hence,

n(N2) = 1.00 g / 28.0134 g mol-1 = 0.0357 mol


n(NaN3) = 2 × n(N2) / 3 = (2 × 0.0357 mol) / 3 = 0.0238 mol


Mass of NaN3 = Moles of NaN3 × Molar mass of NaN3

The molar mass of NaN3 is 65.0099 g mol-1.

Hence,

Mass of NaN3 = 0.0238 mol × 65.0099 g mol-1 = 1.547 g = 1.55 g

Mass of NaN3 = 1.55 g


Answer: 1.55 grams of NaN3 are required.